thabiso claims that if x is a natural number, therefor x2+x will always be an even number.is thabiso correct? explain your answer.


1
Expert's answer
2021-10-04T14:14:39-0400

This is correct

Explanation

"x\\^2 + x\\
=x(x + 1) \\"

Because "x\" is odd number,

so "x + 1" , is even number but any natural number multiplied by even number equals to an even number. Therefore, multiplying "x \times (x + 1)" will give an even number, hence "x\\^2 + x" will always be an even number.